# Exchange control in a MOS double quantum dot made using a 300 mm wafer process

### v2

@@ Abstract
Leveraging the advanced manufacturing capabilities of the semiconductor industry promises to help scale up silicon-based quantum processors by increasing yield, uniformity and integration. Recent studies of quantum dots fabricated on 300 mm wafer metal-oxide-semiconductor (MOS) processes have shown control and readout of individual spin qubits, yet quantum processors require two-qubit interactions to operate. Here, we use a 300 mm wafer MOS process customized for spin qubits and demonstrate coherent control of two electron spins using the spin-spin exchange interaction, forming the basis of an entangling gate such as $\sqrt{\text{SWAP}}$. We observe gate dephasing times of up to $T_{2}^{*}\approx 500$ ns and a gate quality factor of 10. We further extend the coherence by up to an order of magnitude using an echo sequence. For readout, we introduce a dispersive readout technique, the radiofrequency electron cascade, that amplifies the signal while retaining the spin-projective nature of dispersive measurements. Our results demonstrate an industrial grade platform for two-qubit operations, alongside integration with dispersive sensing techniques.
@@
Silicon-based quantum processors have been recently shown to satisfy the requirements for fault-tolerant quantum computing in small systems of up to 2 qubits. Such requirements include qubit initialisation, control, and measurement fidelity above the thresholds required to implement quantum error correcting codes [1, 2, 3, 4, 5, 6]. For electron spin qubits in quantum dots (QDs), further progress has been demonstrated, such as simple instances of error correction in a 3-qubit device [7] and operation of a 6-qubit processor [8] indicating that the technology is ready to be scaled up. To that purpose, it is timely to begin leveraging the manufacturing benefits of the semiconductor industry [9]. This approach promises to produce arrays of qubits with high-yield and reproducibility [10] while facilitating the integration with classical electronics into an integrated quantum computing system [11, 12].
There are several ways to form spin qubits within silicon devices [13, 14], however, those based on metal-oxide-semiconductor (MOS) QDs are of particular interest for industrial manufacturing due to their similarities with silicon field-effect transistors (FETs). As for modern FETs, such QDs can be formed in planar MOS devices [15], or within etched silicon nanowires [16]. The ability to form and operate high quality qubits within silicon MOS structures has been well established [14, 17, 18], motivating recent demonstrations of single qubit performance in MOS devices fabricated using semiconductor manufacturing lines [16, 19]. In this Article, we study electron spins within a silicon double quantum dot (DQD) device fabricated using a 300mm wafer line [10, 20]. We present two key results: First, we introduce a new in-situ dispersive sensing mechanism (termed radiofrequency- electron cascade) with enhanced sensitivity for spin qubit readout. Second, we demonstrate control of an exchange-mediated coherent interaction, which forms the basis for a $\sqrt{\rm SWAP}$ gate between two spin qubits. To achieve the latter, we characterise the angular dependence of the Rashba and Dresselhaus spin-orbit fields to find the optimal in-plane magnetic field direction for operation. We then characterise the exchange interaction as a function of the detuning between the two QDs, and study the effect of an ‘echo’ sequence [21] to mitigate the impact of low-frequency charge noise.
@@
Radiofrequency electron cascade readout
@@
Reading out spin qubits within semiconductor QDs typically involves mapping a spin state of interest onto a charge state of one or more QDs [22, 23], which can then be detected using a variety of charge sensing methods [24]. For example, Pauli spin blockade (PSB) can be used to map the singlet and triplet states of a pair of spin qubits onto two different charge configurations of a DQD (e.g. (1,1) or (0,2)), which are then detected by a single electron transistor [25] or single electron box [26]. In-situ dispersive readout of a DQD combines these into a single step, using PSB to directly distinguish between singlet and triplet states through their difference in AC polarisability [27]. This difference in polarisability is detected by incorporating the DQD into a radiofrequency (rf) tank circuit and measuring changes in the reflected rf signal. However, in-situ dispersive readout has suffered from low sensitivity in planar MOS silicon quantum devices due to the relatively low gate lever arms [28]. To improve the sensitivity of this technique, we introduce a third quantum dot (QD) coupled to a charge reservoir, which acts as an amplifier in measuring the AC-polarisability of the DQD. Instead of measuring the usual single-electron alternating current generated by the cyclic tunneling between the two-spin singlet states of the DQD [24], we leverage the synchronised single-electron AC current at the third dot-reservoir system generated as a consequence of the strong capacitive coupling to the DQD. Such an approach offers the benefit of charge enhancement techniques such as latching [29], dc cascading [30] and spin-polarized single-electron boxes [31] while retaining the non-demolition nature of in-situ dispersive readout methods [32].
@@
Figure 2: |||| Measurement of the spin-orbit interaction. a, Charge stability diagram of the DQD around the (1,1)-(0,2) charge transition with voltage pulse sequence overlaid (shown also as a function of time in panel b, where dashed lines indicate longer durations). c, Energy diagram showing dependence of the two-electron spin states, singlet |$\ket{\text{S}}$ and triplet |$\ket{\text{T}}$ as a function of detuning $\varepsilon$. d, |$\ket{\text{S}}$-|$\ket{\text{T}_{0}}$ oscillations as a function of duration $\tau_{P}$ and detuning at point P $\epsilon_{P}$ in the pulse sequence ($B=250$ mT, $\phi_{B}=235^{\circ}$). e,f, |$\ket{\text{S}}$-|$\ket{\text{T}_{0}}$ oscillation frequency dependence as a function of the in-plane magnetic field orientation $\phi_{B}$, measured at fixed detuning $\varepsilon_{P}=0.926$meV. The fit in f is obtained using the spin-orbit interaction model described in Eq. (2) and incorporates the error in $\sigma_{\text{HF}}$ and $J(\varepsilon_{P})$.
@@
Figure 3: |||| Exchange control in the natSi double quantum device. a, Energy diagram depicting two-electron spin states in the (1,1) detuning regime, with pulse sequence steps overlaid. b, Detuning pulse sequence including initialisation to the $\ket{\downarrow\uparrow}$ state via a semi-adiabatic ramp (orange), followed by a non-adiabatic pulse (J) to near zero-detuning to increase the exchange coupling for duration $\tau_{E}$. c, Exchange driven oscillations between $\ket{\downarrow\uparrow}$ and $\ket{\uparrow\downarrow}$ states, measured rf-phase response proportional to singlet probability, with d, the corresponding Fourier transform. e Ratio between exchange coupling strength $J(\varepsilon)$ and dot-to-dot Zeeman energy difference $\Delta E_{z}$.f Dephasing time $T_{2}^{*}$ extracted from the decay of the exchange oscillations (orange dots) and fit with $T_{2}^{*}=\sqrt{2}\hbar/\delta{\left(h\Omega\right)}$ (purple line) [21, 33]. g Gate quality factor.
@@
We use a device based on planar silicon MOS technology with an overlapping gate design (see Fig 1(a)). Quantum dots Q1 and Q2 form a DQD which we tune to hold two electrons, while Q2 is capacitively coupled to a multi-electron quantum dot (QME) that can exchange electrons with a charge reservoir. To measure the charge state of the system, we connect a superconducting spiral inductor to the reservoir forming an $LC$ resonator (see Fig. 1(b)). At voltages where the charge in QME is bistable, cyclic tunneling generated by the small rf signal supplied to the resonator produces a change in capacitance that can be detected as a change in the phase response, $\Delta\Phi$, using homodyne techniques [24]. Lines in gate voltage space showing QME charge bistability are shifted when intersecting charge transitions of the DQD, as shown in Fig. 1(c). Such shifts form the basis of dispersive charge sensing measurements [34, 26, 35] which we do not exploit here. Instead, we focus on the directly observable AC signal in the region of gate voltage space where charge transitions may occur between Q1 and Q2. We ascribe this signal to a two-electron charge cascade effect driven by the rf excitation, which we explain using the diagrams in Fig. 1(d-e). Consider an rf cycle in which the system starts in the occupation configuration ($\text{N}_{\text{Q}_{1}}$,$\text{N}_{\text{Q}_{2}}$,$_{\text{Q}_{\text{ME}}})=(1,1,\text{N})$ N ). Due to the strong capacitive coupling between Q2 and QME, the rf excitation that drives the DQD from the $(1,1)$ to $(0,2)$ state synchronously forces an electron out of QME in a cascaded manner, leading to $(0,2,\text{N}-1)$. The second half of the rf cycle then reverses the process. Overall, the rf cascade measures the polarizability of the DQD system, as for in-situ dispersive readout measurements [28, 36], but with the substantial advantage that the induced charge can be much greater, resulting in a dispersive measurement with greater sensitivity (see App. APPENDIX A: ELECTROSTATIC REQUIREMENTS FOR RF-DRIVEN ELECTRON CASCADE READOUTA for more discussion on the requirements for rf electron cascade readout). Specifically, we find the voltage signal is amplified in the cascade approach compared to direct dispersive readout, by a factor
$A=1+\frac{1-\alpha_{\text{r,ME}}}{\alpha_{\text{r,2}}-\alpha_{\text{r,1}}}>1,$ (1)
where $\alpha_{{\rm r},j}$ represents the gate lever arm between the reservoir and QD $j$. Particularly, the sensor detects not only the interdot gate polarization charge $(\alpha_{\text{r,1}}-\alpha_{\text{r,2}})e$ but also the cascaded charge collected at the reservoir, $(1-\alpha_{\text{r,ME}})e$, (see App. APPENDIX B: AMPLIFICATION FACTORB). We highlight that, as opposed to other charge enhancement techniques [29, 31, 30], the rf electron cascade retains the non-demolition nature of in-situ dispersive readout measurements since the DQD system remains in an eigenstate after a measurement is performed [37, 38].
@@
We use the rf-cascade to distinguish the singlet and triplet states of the DQD via PSB. The signature of PSB can be observed by measuring the asymmetric disappearance of the interdot charge transitions as a function of increasing applied magnetic field [27], as shown in the inset to Fig. 1(c). At low magnetic field, the system is free to oscillate between singlet states (|$\ket{\text{S}(1,1)}\leftrightarrow\ket{\text{S}(0,2)}$) due to the action of the rf drive, yielding a signal in the rf response. At high fields, however, the separated triplet |$\ket{\text{T}_{-}(1,1)}$ state becomes the ground state for all relevant values of detuning, preventing a charge transition and resulting in the disappearance of the signal, initially for the region of the transition closer to the $(1,1)$ charge configuration. A quantum capacitance-based simulation of the data in the inset suggests an inter-dot tunnel coupling of $t_{c}=2.4$ GHz and electron temperature $T_{e}=50$ mK (see App. APPENDIX C: MAGNETOSPECTROSCOPY SIMULATIONC). We note that the rf excitation is continuously applied for all measurements in this Article.
@@
Having established a method to distinguish between singlet and triplet spin states, our goal is to prepare and coherently control spin states of the DQD through voltage pulses along the detuning axis, $\epsilon$. Such pulses bring the DQD: i) from the (2,0) charge configuration in which a singlet is prepared; ii) into the (1,1) region where the electron spins are spatially separated between QDs and may evolve; and iii) back to an intermediate point where they can be measured (see Fig. 2(a-c)). Deep in the (1,1) region, the spin basis states are predominantly $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$. Under an adiabatic ramp to $\epsilon=0$ for readout, these two states map onto the |$\ket{\text{T}_{0}(1,1)}$ and |$\ket{\text{S}(0,2)}$, respectively. The basis states are separated in energy by $h\Omega=\sqrt{J(\varepsilon)^{2}+\Delta E_{z}^{2}}$, where we include the kinetic exchange interaction $J(\varepsilon)$ and the Zeeman energy difference between electrons in each dot $\Delta E_{z}$. The spin detuning $\Delta E_{z}=\Delta g\mu_{B}B+g\mu_{B}\Delta B_{\text{HF}}$ (where $\mu_{B}$ is the Bohr magneton, $h$ is Planck’s constant) contains two main contributions: (i) the difference in g-factor between QDs $\Delta g=\left|g_{2}-g_{1}\right|$ arising from variations in the spin-orbit interaction (SOI) present near the Si/SiO2 interface  [39, 40, 18]; and (ii) the difference in the effective 29Si nuclear magnetic field experienced by each QD, $\Delta B_{\text{HF}}$. The random fluctuations in the effective magnetic field experienced by each electron in the DQD can be described by a normal distribution with mean of 0 (given the negligible spin polarisation) and standard deviation $\sigma_{\text{HF}}=0.13\pm 0.02$ mT, as we shall see later. This value corresponds to a hyperfine energy strength of $3.4\pm 0.4$ neV, which aligns well with other reports in natural silicon [41, 33, 42].
@@
We implement exchange control using the sequence depicted in Fig. 3(a,b), where the $\ket{\downarrow\uparrow}$ state is initialised via a ramp from $\varepsilon=0$ into the (1,1) configuration that is adiabatic with respect to $E_{z}$ [46]. A fast non-adiabatic pulse towards zero detuning increases the exchange coupling, driving oscillations between the $\ket{\downarrow\uparrow}$ and $\ket{\uparrow\downarrow}$ states at frequency $\Omega(\varepsilon)$, as observed in Fig. 3(c). The final state after some evolution time $\tau_{J}$ is projected to |$\ket{\text{S}}$ or |$\ket{\text{T}_{0}}$ for readout. The Fourier transform of the exchange oscillations (see Fig. 3(d)) reveals a single peak of increasing frequency as the detuning is reduced, indicating the purity of the oscillations and the enhanced exchange strength at lower detuning.
To quantify the properties of these rotations, we combine the results of the exchange oscillations in Fig. 3(c) and the |$\ket{\text{S}}$-|$\ket{\text{T}_{0}}$ oscillations in Fig. 2(d), to extract the ratio $J/\Delta E_{z}$ and the intrinsic coherence time $T_{2}^{*}$ over a wide range of detunings, see Fig. 3(e,f). We extract $T_{2}^{*}$ by fitting the oscillations at each detuning point with a Gaussian decay envelope of the form exp$\left[(-\tau/T_{2}^{*})^{2}\right]$, and then obtain $\Delta E_{z}=9.6\pm 1.2~{}$MHz from the fit to the expression
$\frac{1}{T_{2}*}=\frac{1}{\sqrt{2}\hbar}\sqrt{\left(\frac{J}{h\Omega}\frac{dJ}{d\varepsilon}\delta\varepsilon_{\text{rms}}\right)^{2}+\left(\frac{\Delta E_{z}}{h\Omega}\delta\Delta E_{z,\text{rms}}\right)^{2}},$ (3)
where $\delta\varepsilon_{\text{rms}}$ and $\delta\Delta E_{z,\text{rms}}$ refer to the root mean square of the fluctuations in $\varepsilon$ and $\Delta E_{z}$ [21, 33, 43].

### v3

# Radio-frequency cascade readout and coherent exchange control of spin qubits fabricated using a 300 mm wafer process

@@ Abstract
Leveraging advanced semiconductor manufacturing promises significant potential for scaling up silicon-based quantum processors. In this context, control and readout of individual spin qubits have been shown on devices manufactured on 300 mm wafer metal-oxide-semiconductor (MOS) processes, yet quantum processors require two-qubit interactions to operate. Here, we use a 300 mm natural silicon MOS process customized for spin qubits and demonstrate coherent control of two electron spins using the spin-spin exchange interaction, which forms the basis for entangling gates such as $\sqrt{\text{SWAP}}$. We observe gate dephasing times of up to $T_{2}^{*}\approx 500$ ns and a gate quality factor of 10. For readout, we introduce a novel dispersive readout technique, the radio-frequency electron cascade, that simplifies the qubits unit cell while providing high gain. We achieve a 74% singlet-triplet readout fidelity in 8 $\upmu$s integration time, the highest-performing dispersive readout demonstration on a planar MOS process. The combination of sensitive dispersive readout with industrial-grade manufacturing marks a crucial step towards the large-scale integration of silicon quantum processors.
@@
Silicon-based quantum processors have recently demonstrated high-fidelity qubit initialisation [1, 2], measurement [3, 4, 5], and single-[6, 7, 8] and two-qubit control [9, 10, 11, 12, 13, 14] in small-scale devices of up to two qubits, with fidelities exceeding the $99\%$ threshold required to implement quantum error correction [15]. Further progress has been achieved using electron spin qubits in quantum dots (QDs). In particular, simple instances of quantum error correction have been demonstrated in a 3-qubit array [2], and a 6-qubit processor has been successfully operated [16]. As the technology matures, it is crucial to take advantage of the manufacturing capabilities of the semiconductor industry to scale up to larger arrays of qubits. This approach promises to facilitate integration with classical electronics for control and readout, paving the way for an integrated quantum computing system [17, 18, 19].
Spin qubits based on silicon metal-oxide-semiconductor (MOS) technology  [20, 21] and Si/SiGe heterostructures [22, 23, 16, 11, 24] are of particular interest for industrial manufacturing. Specifically, the MOS approach shares similarities with modern silicon field-effect transistor (FET) manufacturing, enabling the formation of MOS QDs in both planar devices [13, 20, 21, 25, 26, 27], and etched silicon structures, such as nanowires [28, 29, 30] or finFETs [6, 31]. While the latter constrains the qubit topology to 2xN bilinear QD arrays, the former offers easier scalability towards two-dimensional QD arrays [32], which are essential for implementing quantum error correcting codes such as the surface code. This potential for scaling has motivated the translation of the well-established operation of high-quality qubits within MOS academic devices [8, 13, 25, 20, 26], to demonstrations of single-qubit performance in MOS devices fabricated using semiconductor manufacturing lines [28, 31]. Yet, quantum processors require two-qubit interactions to operate.
For further scale-up, methods that simplify the readout infrastructure must be developed. The current standard for sensing, the radio-frequency single-electron transistor, offers high-fidelity readout [3, 5] at the cost of occupying substantial space on the qubit chip, which limits qubit connectivity. To address this limitation, efforts have focused on integrating fast and compact radio-frequency measurement techniques [33, 34, 35], which substantially reduce the readout footprint. However, despite recent advances in other areas, dispersive readout metrics have seen little progress in planar MOS in recent years [34].
@@
In this Article, we study electron spins within a silicon double quantum dot (DQD) device fabricated using a 300 mm wafer line [36, 37] to address the two aforementioned outstanding challenges: First, we introduce a new in situ dispersive sensing mechanism (termed radiof-requency- electron cascade) which offers enhanced sensitivity for spin qubit readout. We achieve a minimum integration time of 13 $\upmu$s and using a physical model, we calculate 74% singlet-triplet readout fidelity in an 8 $\upmu$s integration time. Second, we demonstrate control of an exchange-mediated coherent interaction, which forms the basis for a $\sqrt{\rm SWAP}$ gate between two spin qubits.
Radio-frequency electron cascade readout
@@
Reading out spin qubits within semiconductor QDs typically involves mapping a spin state of interest onto a charge state of one or more QDs [38, 39], which can then be detected using a variety of charge sensing methods [40]. For example, Pauli spin blockade (PSB) can be used to map the singlet and triplet states of a pair of spin qubits onto two different charge configurations of a DQD (e.g. (1,1) or (0,2)), which are then detected by a single electron transistor [41] or single electron box [4]. In-situ dispersive readout of a DQD combines these into a single step, using PSB to directly distinguish between singlet and triplet states through their difference in AC polarisability [42]. This difference in polarisability is detected by incorporating the DQD into a radio-frequency (rf) tank circuit and measuring changes in the reflected rf signal. However, in-situ dispersive readout has suffered from low sensitivity in planar MOS silicon quantum devices due to the relatively low gate lever arms [34]. To improve the sensitivity of this technique, we introduce a third quantum dot (QD) coupled to a charge reservoir, which acts as an amplifier in measuring the AC-polarisability of the DQD. Instead of measuring the usual single-electron alternating current generated by the cyclic tunneling between the two-spin singlet states of the DQD [40], we leverage the synchronised single-electron AC current at the third dot-reservoir system generated as a consequence of the strong capacitive coupling to the DQD. Our approach offers the benefit of charge enhancement techniques such as latching [43], DC cascading [44] and spin-polarized single-electron boxes [45] while retaining the non-demolition nature of in-situ dispersive readout methods [46].
Fig. 2: |||| Radio-frequency cascade readout fidelity. a Charge stability diagram of the DQD around the (1,1)-(0,2) interdot charge transition (ICT), with a voltage pulse sequence and detuning axis $\varepsilon$ overlaid b, Reflected resonator voltage $V_{\rm rf}$ as a function of detuning across the ICT around $\varepsilon=0$, fit to eq. (5) (dashed orange line), using tunnel coupling $t_{c}$. c Energy diagram showing dependence of the two-electron spin states, singlet |$\ket{\text{S}}$ and triplet |$\ket{\text{T}}$ as a function of $\varepsilon$. The pulse sequence depicts the initialisation (I) of the |$\ket{{\rm S(0,2)}}$ via an adiabatic ramp from the $(0,1)$ empty (E) state, followed by a non-adiabatic pulse to $\varepsilon=0$ for measurement (M). d, $V_{\rm rf}$ as a function of wait time $\tau_{\rm M}$ at $\varepsilon=0$ before measurement, following the pulse sequence depicted in panels a and c. The fitted line indicates a |$\ket{\text{S}}$ to |$\ket{{\rm T_{-}(1,1)}}$ relaxation time $T_{1}=24~{}\upmu$s. e, Calculated readout infidelity $1-\mathcal{F}_{r}$ as a function of integration time $\tau_{\text{int}}$ as described by Eq. (2). Solid lines depict relaxation time $T_{1}$ limited fidelity, whereas dashed lines depict $T_{1}=\infty$. Parameters used from left to right include: $\tau_{\text{min}}=\left\{94~{}\text{ns},13~{}\upmu\text{s},"\right\}$, $T_{1}=\left\{24~{}\upmu\text{s},",4.5~{}\text{ms}\right\}$.
@@
Fig. 3: |||| Measurement of the spin-orbit interaction. a, Schematic of the DQD charge stability diagram around the (1,1)-(0,2) interdot charge transition with voltage pulse sequence overlaid (shown also as a function of time in panel b, where dashed lines indicate longer durations). c, Energy diagram showing dependence of the two-electron spin states, singlet |$\ket{\text{S}}$ and triplet |$\ket{\text{T}}$ as a function of detuning $\varepsilon$. d, |$\ket{\text{S}}$-|$\ket{\text{T}_{0}}$ oscillations as a function of duration $\tau_{P}$ and detuning at point P $\epsilon_{P}$ in the pulse sequence ($B=250$ mT, $\phi_{B}=235^{\circ}$). e-f, |$\ket{\text{S}}$-|$\ket{\text{T}_{0}}$ oscillation frequency dependence as a function of the in-plane magnetic field orientation $\phi_{B}$, measured at fixed detuning $\varepsilon_{P}=0.926$meV. The fit in f is obtained using the spin-orbit interaction model described in Eq. (3) and incorporates the error in $\sigma_{\text{HF}}$ and $J(\varepsilon_{P})$.
@@
Fig. 4: |||| Exchange control in the natSi double quantum dot device. a, Energy diagram depicting two-electron spin states in the (1,1) detuning regime, with pulse sequence steps overlaid. b, Detuning pulse sequence including initialisation to the $\ket{\downarrow\uparrow}$ state via a semi-adiabatic ramp (orange), followed by a non-adiabatic pulse (J) to near zero-detuning to increase the exchange coupling for duration $\tau_{E}$. c, Exchange driven oscillations between $\ket{\downarrow\uparrow}$ and $\ket{\uparrow\downarrow}$ states, measured rf-phase response proportional to singlet probability, with d, the corresponding Fourier transform. e Ratio between exchange coupling strength $J(\varepsilon)$ and dot-to-dot Zeeman energy difference $\Delta E_{z}$. f Dephasing time $T_{2}^{*}$ extracted from the decay of the exchange oscillations (orange dots) and fit with $T_{2}^{*}=\sqrt{2}\hbar/\delta{\left(h\Omega\right)}$ (purple line) [47, 48]. g Qubit quality factor. Error bars indicate standard deviation of each data point.
@@
We use a device based on planar silicon MOS technology with an overlapping gate design [37] (see Fig 1(a)). Quantum dots Q1 and Q2 form a DQD which we tune to hold two electrons, while Q2 is capacitively coupled to a multi-electron quantum dot (QME) that can exchange electrons with a charge reservoir. To measure the charge state of the system, we connect a superconducting spiral inductor to the reservoir, forming a $LC$ resonator (see Fig. 1(b)). At voltages where the charge in QME is bistable, cyclic tunneling generated by the small rf signal supplied to the resonator produces a change in capacitance that can be detected as a change in the phase response, $\Delta\Phi$, using homodyne techniques [40]. Lines in gate voltage space showing QME charge bistability are shifted when intersecting charge transitions of the DQD, as shown in Fig. 1(c). Such shifts form the basis of dispersive charge sensing measurements [49, 4, 50] which we do not exploit here. Instead, we focus on the directly observable AC signal in the region of gate voltage space where charge transitions may occur between Q1 and Q2. We ascribe this signal to a two-electron charge cascade effect driven by the rf excitation, which we explain using the diagrams in Fig. 1(d-e). Consider an rf cycle in which the system starts in the occupation configuration ($\text{N}_{\text{Q}_{1}}$,$\text{N}_{\text{Q}_{2}}$,$_{\text{Q}_{\text{ME}}})=(1,1,\text{N})$ N ). Due to the strong capacitive coupling between Q2 and QME, the rf excitation that drives the DQD from the $(1,1)$ to $(0,2)$ state synchronously forces an electron out of QME in a cascaded manner, leading to $(0,2,\text{N}-1)$. The second half of the rf cycle then reverses the process. Overall, the rf cascade measures the polarizability of the DQD system, as for in-situ dispersive readout measurements [34, 35], but with the substantial advantage that the induced charge can be much greater, resulting in a dispersive measurement with greater sensitivity (see Supplementary Section S1 for more discussion on the requirements for rf electron cascade readout). Specifically, we find the power signal is amplified in the cascade approach compared to direct dispersive readout, by a factor
$A=\left(1+\frac{1-\alpha_{\text{R,ME}}}{\alpha_{\text{R,2}}-\alpha_{\text{R,1}}}\right)^{2}>1,$ (1)
where $\alpha_{{\rm R},j}$ represents the gate lever arm between the reservoir and QD $j$. In particular, the sensor detects not only the interdot gate polarization charge $(\alpha_{\text{R,1}}-\alpha_{\text{R,2}})e$ but also the cascaded charge collected at the reservoir, $(1-\alpha_{\text{R,ME}})e$, (see Supplementary Section S2 for derivation). By comparing the measured signal-to-noise ratio (SNR) with and without the cascade effect, we find a lower bound for the power amplification factor, $A\geq 1010\pm 40$ $(+36~{}\text{\rm dB})$. The cascade SNR is extracted from the interdot charge transition fit shown in Fig. 2(a-b) (see eq. (5-6) in Methods). For the case where the cascade is absent, the SNR is assumed to be 0.5, as there is no observable signal, as shown in Supplementary Fig. S1(a) and (e). From the cascade SNR we extract a minimum integration time of $\tau_{\rm min}=13.0\pm 0.5~{}\upmu{\rm s}$ (see eq. (7) in Methods), representing an improvement of over two orders-of-magnitude on prior planar MOS demonstrations of in-situ dispersive readout [34].
@@
We use the rf-cascade to distinguish between the singlet and triplet states of the DQD via PSB. The signature of PSB can be observed by measuring the asymmetric disappearance of the interdot charge transitions as a function of increasing the applied magnetic field [42], as shown in the inset of Fig. 1(c). At low magnetic field $(B\leq 200~{}\text{\rm mT})$, the system is free to oscillate between singlet states (|$\ket{\text{S}(1,1)}\leftrightarrow\ket{\text{S}(0,2)}$) due to the action of the rf drive, yielding a signal in the rf response. However, at higher fields $(B\geq 200~{}\text{\rm mT})$, the polarised triplet |$\ket{\text{T}_{-}(1,1)}$ state becomes the ground state for $\varepsilon\geq 0$ (as shown in Fig. 2(c)), preventing a charge transition and resulting in the disappearance of the signal, initially for the region of the transition closer to the $(1,1)$ charge configuration. A quantum capacitance-based simulation of the data in the inset suggests an inter-dot tunnel coupling of $t_{c}=2.4$ GHz and electron temperature $T_{e}=50$ mK (see Supplementary Section S3). Figure 2(d) shows the decay from |$\ket{\text{S}}$ to |$\ket{\text{T}_{-}(1,1)}$ at $\varepsilon=0$ for applied magnetic field $B=250~{}$mT. The corresponding relaxation time is $T_{1}=24\pm 7~{}\upmu$s, which is consistent with a study of a similar device, where a relaxation time of $T_{1}=10~{}\upmu{\rm s}$ was observed near zero detuning for a comparable tunnel coupling ($t_{c}\geq 1.9~{}$GHz) [51]. In that study, $T_{1}$ was found to vary exponentially with $t_{c}$, with relaxation times exceeding $200~{}\upmu$s for $t_{c}\leq 1~{}$GHz.
We assess the performance of the radio-frequency driven electron cascade by calculating the readout fidelity, $\mathcal{F}_{r}$ as a function of integration time, using the expression [40, 52],
$\mathcal{F}_{r}=\frac{1}{2}\left[1+\text{erf}\left(\sqrt{\frac{\tau_{\text{int}}}{2\tau_{\text{min}}}}\right)\exp\left(\frac{\tau_{\text{int}}}{2T_{1}}\right)\right].$ (2)
We calculate a readout fidelity of $\mathcal{F}_{r}=74\begin{subarray}{c}+1\\ -2\end{subarray}\%$ based on the experimentally obtained $\tau_{\rm min}=13~{}\upmu$s and integration time $\tau_{\text{int}}=8~{}\upmu$s. The corresponding readout infidelity $1-\mathcal{F}_{r}$ is depicted in Fig. 2(c) (dark purple line). This fidelity is comparable to prior in-situ readout demonstrations, but is achieved with integration times that are 1-2 orders of magnitude faster than those reported in previous silicon planar MOS and implanted donor systems [34, 33]. Our performance approaches that of integrated on-chip resonators in Si/SiGe [35], but without the added fabrication complexity of on-chip resonator integration. To achieve further improvements in fidelity, two strategies are indicated in Fig. 2(c): ($i$) reducing the minimum integration time to 949494~{}94ns, to achieve a fidelity of $99\%$ (light purple line), by enhancing the SNR through resonator optimisation or in-situ matching techniques [40, 53]; and ($ii$) by extending the relaxation time of the system [51] to that reported in the state-of-the-art MOS [34] (orange line), resulting in an estimated fidelity of $99.2\%$.
We highlight that, unlike other charge enhancement techniques [43, 45, 44], the rf electron cascade retains the non-demolition nature of in-situ dispersive readout measurements since the DQD system remains in an eigenstate after a measurement is performed [54, 1]. We note that the rf excitation is continuously applied for all measurements in this Article.
@@
Having established a method to distinguish between singlet and triplet spin states, our goal is to prepare and coherently control spin states of the DQD through voltage pulses along the detuning axis, $\epsilon$. Such pulses bring the DQD: i) from the (0,2) charge configuration in which a singlet is prepared; ii) into the (1,1) region where the electron spins are spatially separated between QDs and may evolve; and iii) back to an intermediate point where they can be measured (see Fig. 3(a-c)). Deep in the (1,1) region, the spin basis states are predominantly $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$. Under an adiabatic ramp to $\epsilon=0$ for readout, these two states map onto the |$\ket{\text{T}_{0}(1,1)}$ and |$\ket{\text{S}(0,2)}$, respectively. The basis states are separated in energy by $h\Omega=\sqrt{J(\varepsilon)^{2}+\Delta E_{z}^{2}}$, where we include the kinetic exchange interaction $J(\varepsilon)$ and the Zeeman energy difference between electrons in each dot $\Delta E_{z}$. The spin detuning $\Delta E_{z}=\Delta g\mu_{B}B+g\mu_{B}\Delta B_{\text{HF}}$ (where $\mu_{B}$ is the Bohr magneton, $h$ is Planck’s constant) contains two main contributions: (i) the difference in g-factor between QDs $\Delta g=\left|g_{2}-g_{1}\right|$ arising from variations in the spin-orbit interaction (SOI) present near the Si/SiO2 interface  [55, 56, 26]; and (ii) the difference in the effective 29Si nuclear magnetic field experienced by each QD, $\Delta B_{\text{HF}}$. The random fluctuations in the effective magnetic field experienced by each electron in the DQD can be described by a normal distribution with mean of 0 (given the negligible spin polarisation) and standard deviation $\sigma_{\text{HF}}=30\pm 4~{}\upmu$T, as we shall see later. This value corresponds to a hyperfine energy strength of $3.4\pm 0.4$ neV, which aligns well with other reports in natural silicon [22, 48, 57].
@@
We implement exchange control using the sequence depicted in Fig. 4(a,b), where the $\ket{\downarrow\uparrow}$ state is initialised via a ramp from $\varepsilon=0$ into the (1,1) configuration that is adiabatic with respect to $E_{z}$ [60, 22]. A fast non-adiabatic pulse towards zero detuning increases the exchange coupling, driving oscillations between the $\ket{\downarrow\uparrow}$ and $\ket{\uparrow\downarrow}$ states at frequency $\Omega(\varepsilon)$, as observed in Fig. 4(c). The final state after some evolution time $\tau_{J}$ is projected to |$\ket{\text{S}}$ or |$\ket{\text{T}_{0}}$ for readout. The Fourier transform of the exchange oscillations (see Fig. 4(d)) reveals a single peak of increasing frequency as the detuning is reduced, indicating the purity of the oscillations and the enhanced exchange strength at lower detuning. From this, we observe that the exchange coupling is tunable over a range of $J=5-122~{}$MHz.
To quantify the properties of these rotations, we combine the results of the exchange oscillations in Fig. 4(c) and the |$\ket{\text{S}}$-|$\ket{\text{T}_{0}}$ oscillations in Fig. 3(d), to extract the ratio $J/\Delta E_{z}$ and the intrinsic coherence time $T_{2}^{*}$ over a wide range of detunings, see Fig. 4(e,f). We extract $T_{2}^{*}$ by fitting the oscillations at each detuning point with a Gaussian decay envelope of the form exp$\left[-(\tau/T_{2}^{*})^{2}\right]$, and then obtain $\Delta E_{z}=9.6\pm 1.2~{}$MHz from the fit to the expression
$\frac{1}{T_{2}^{*}}=\frac{1}{\sqrt{2}\hbar}\sqrt{\left(\frac{J}{h\Omega}\frac{dJ}{d\varepsilon}\delta\varepsilon_{\text{rms}}\right)^{2}+\left(\frac{\Delta E_{z}}{h\Omega}\delta\Delta E_{z,\text{rms}}\right)^{2}},$ (4)
where $\delta\varepsilon_{\text{rms}}$ and $\delta\Delta E_{z,\text{rms}}$ refer to the root mean square of the fluctuations in $\varepsilon$ and $\Delta E_{z}$ [47, 48].